# HiCMAE: Hierarchical Contrastive Masked Autoencoder for Self-Supervised Audio-Visual Emotion Recognition

### v1

@@ 4.3.3 Results on MSP-IMPROV, RAVDESS, and IEMOCAP
The audio-visual results on MSP-IMPROV busso2016msp are shown in Table 7. We observe that, when compared with the state-of-the-art AW-HuBERT, HiCMAE-B achieves similar UAR but shows large improvement (+3.15%) in terms of WAR. Two smaller versions of HiCMAE have lower UAR. However, their WAR is still higher than AW-HuBERT.
@@
LR+eGeMAPS keesing2023emotion ; eyben2015geneva A - - 50.30 -
LR+wav2vec keesing2023emotion ; baevski2020wav2vec A - - 68.80 -
Wav2Vec2.0 baevski2020wav2vec A 95 18 73.44 74.38
HuBERT hsu2021hubert A 95 18 74.15 74.37
WavLM-Plus chen2022wavlm A 95 18 75.28 75.36
@@
VO-LSTM ghaleb2019multimodal $\times$ V - - - 60.50
3D ResNeXt-50 su2020msaf $\times$ V 26 - - 62.99
SVFAP sun2023svfap V 78 44 75.15 75.01
MAE-DFER sun2023mae V 85 50 75.91 75.56
@@
AV-LSTM ghaleb2019multimodal $\times$ A+V - - - 65.80
AV-Gating ghaleb2019multimodal $\times$ A+V - - - 67.70
MCBP fukui2016multimodal $\times$ A+V 51 - - 71.32
MMTM joze2020mmtm $\times$ A+V 32 - - 73.12
MSAF su2020msaf $\times$ A+V 26 - - 74.86
ERANNs verbitskiy2022eranns $\times$ A+V - - - 74.80
CFN-SR fu2021cross $\times$ A+V 26 - - 75.76
MATER ghaleb2020multimodal $\times$ A+V - - - 76.30
MulT tsai2019multimodal $\times$ A+V - - - 76.60
AVT chumachenko2022self $\times$ A+V - - - 79.20
@@
Attn. Pooling sadok2023vector
@@
FBANK tseng2023av $\times$ A - - - 51.52
AV-HuBERT shi2022learning A 90 - - 58.54
RepLAI mittal2022learning A 5 - - 57.53
AVBERT lee2021parameter A 10 - - 60.94
MAViL lee2021parameter A 86 - - 59.46
Wav2vec 2.0 baevski2020wav2vec A 95 18 69.88 67.32
HuBERT hsu2021hubert A 95 18 68.33 66.34
WavLM-Plus chen2022wavlm A 95 18 68.64 67.12
@@
HoG dalal2005histograms $\times$ V - - - 35.83
AV-HuBERT shi2022learning V 103 - - 26.59
RepLAI mittal2022learning V 15 - - 40.72
AVBERT lee2021parameter V 37 - - 45.80
MAViL lee2021parameter V 87 - - 43.03
@@
AV-HuBERT shi2022learning A+V 103 - - 46.45
AVBERT lee2021parameter A+V 43 - - 61.87
MAViL lee2021parameter A+V 187 - - 54.94
@@
In Table 8, we compare the proposed method with state-of-the-art methods on RAVDESS livingstone2018ryerson . It can be seen that HiCMAE brings large gains over the previous best method. In specific, our HiCMAE-B achieves a gain of +3.19% WAR over VQ-MAE-AV+Query2Emo sadok2023vector . HiCMAE-T still outperforms it by +1.31% WAR, while using 33% fewer parameters. The unimodal results are less satisfactory on this dataset, probably due to the much smaller model size.
The performance comparison on IEMOCAP busso2008iemocap are presented in Table 9. Most baseline results are from AV-SUPERB tseng2023av . We find that HiCMAE achieves substantial improvement over state-of-the-art generic audio-visual representation learner (e.g., MAViL huang2023mavil and AVBERT lee2021parameter ) in the audio-visual setting. For example, our HiCMAE-S surpasses AVBERT by +5.62% WAR while having a similar model size. HiCMAE-T outperforms MAViL which is also built upon masked autoencoders by +11.68% WAR while being 9$\times$ smaller, indicating the importance of reducing domain shift and the benefit of hierarchical feature learning to improve the quality of the learned representations. When evaluating visual-only performance, our method also demonstrates great success. However, the audio-only results are inferior to state-of-the-art large pre-trained speech models, leaving much room for future performance improvement.
@@ 4.4.1 Datasets
Werewolf-XL zhang2021werewolf is an audio-visual database for studying spontaneous emotions during competitive group interactions in Werewolf games. It contains a total of about 15 hours of audio-visual recordings. In this paper, we use 14,632 speakers’ samples with dimensional annotations (i.e., arousal, valence, and dominance) and conduct subject-independent 5-fold cross-validation for model evaluation.
AVCAffe sarkar2022avcaffe is a large-scale audio-visual affect dataset simulating remote work scenarios. It consists of a total of 108 hours of videos (more than 58,000 video clips) along with self-reported labels for cognitive load and affect (i.e., arousal, and valence). Note that the arousal and valence scores are given on a scale of 1-4 and we follow the original paper to formulate their prediction as a classification task instead of a regression one. This dataset provides official split (86 subjects for training and 20 subjects for test) for model evaluation.
@@
eGeMAPS eyben2015geneva $\times$ A 23.45 28.16 8.08 6.86 31.15 34.03
VGGish hershey2017cnn $\times$ A 22.88 30.50 5.69 3.91 29.59 31.57
@@
HOG dalal2005histograms $\times$ V 20.82 14.43 52.54 34.56 24.76 16.90
VGGFace parkhi2015deep $\times$ V 7.24 4.61 62.96 60.38 14.30 8.20
SVFAP sun2023svfap V 23.51 18.96 67.11 64.27 34.61 29.69
@@
Zhang et al. zhang2021werewolf $\times$ A+V 16.41 27.70 63.14 62.34 35.40 38.40
@@ 4.4.2 Results
In Table 10, we compare HiCMAE with state-of-the-art methods on Werewolf-XL zhang2021werewolf . It can be seen that our method outperforms baselines by large margins in terms of Pearson correlation coefficient (PCC) and concordance correlation coefficient (CCC). In the audio-visual setting, for example, HiCAME-B surpasses the previous best by +17% PCC and +4% CCC in arousal and +6% PCC and +4% CCC in valence. In the unimodal setting, HiCMAE also achieves competitive or even better results. Specifically, for the video modality, HiCMAE-B has similar performance with the state-of-the-art SVFAP while having significantly fewer parameters (32M versus 78M) and FLOPs (32G versus 44G). For the audio modality, HiCMAE brings large performance gains over the best-performing baselines on two evaluation metrics in three emotion dimensions except for CCC in arousal.
The audio-visual performance comparison on AVCAffe sarkar2022avcaffe is presented in Table 11. We find that HiCMAE-B outperforms the previous best results by +2.68% weighted F1-score in arousal and +3.50% weighted F1-score in valence. As for the smaller model HiCMAE-S, although we see some performance drop, it still beats the baselines moderately in two dimensions while having similar parameters with them. Finally, for the smallest version HiCMAE-T, it typically lags behind the previous supervised methods, probably due to its too small model capacity.
@@ 4.5 Ablation Studies
In this section, we conduct in-depth ablation studies to investigate several key design factors in HiCMAE. By default, we present the results of HiCMAE-B on two representative datasets, i.e., MAFW and CREMA-D (6-class).
Table 12: Ablation study on three modules for hierarchical feature learning. HSP: hierarchical skip connections. HCMCL: hierarchical cross-modal contrastive learning. HFF: hierarchical feature fusion.
@@ 4.5.1 Three-pronged Strategy for Hierarchical Feature Learning
Table 13: Ablation study on loss weights.
@@ 4.5.4 Information Flow in Cross-modal Fusion Encoder
Figure 4: Four different types of information flow in cross-modal fusion encoder.
@@
raw input 42.30 55.86 84.47 84.43
video first 42.74 56.06 84.81 84.78
audio first 42.44 56.00 84.60 84.57
@@
We then investigate the effect of different types of information flow in the cross-modal fusion encoder. We develop three variants of the default information flow in Eq. (5-6) and show their differences in Fig. 4. Specifically, for the raw input variant, tokens of one modality in each fusion layer always attend to the raw input tokens of the other modality tsai2019multimodal , instead of updated tokens from the last layer. For the video first variant, video tokens first update themselves via audio information from the last fusion layer and then audio tokens attend to the updated video tokens. The audio first variant is just the reverse of the video first variant. The ablation results are presented in Table 15. We observe that the model performance is not sensitive to different types of information flow in the cross-modal fusion encoder. Besides, in general, the default information flow works best, followed by the video first and audio first variants, and finally the raw input variant.
@@ 4.6.1 Masked Audio-Visual Reconstruction
We first show the reconstruction ability of HiCMAE under different masking rates. We randomly select video clips of four celebrities from the unseen test set of VoxCeleb2 for evaluation. The results of masked audio-visual reconstruction are shown in Fig. 7. For each video clip, we display the original audio spectrogram and 8 facial frame images in the first row, the masked input in medium level (60% for audio and 75% for video) along with the corresponding reconstructed data in the second and third row, and the highly masked input (80% for audio and 90% for video) along with the corresponding reconstructed data in the last two rows. From the figure, we observe that, although some fine-grained details are lost, the global structures of audio spectrogram (e.g., harmonics) and visual frames (e.g., smile) are well recovered under both medium and high masking rates. These satisfactory reconstruction results indicate that HiCMAE is capable of fully exploiting limited intra-modal and cross-modal contextual information to infer the missing audio-visual data. We believe this capability learned during self-supervised pre-training has laid the foundation for its superior downstream fine-tuning performance.
@@ 4.6.2 Embedding Space
In this part, we utilize t-SNE van2008visualizing to visualize the learned feature embedding space of HiCMAE on CREMA-D (6-class). To demonstrate the effect of audio-visual fusion, we present both unimodal and multimodal embedding space. The results are shown in Fig. 8. Each row presents unimodal or audio-visual embedding space and each column denotes one of five folds on this dataset. From the figure, we find that both audio-only and video-only models have learned good embedding space for distinguishing different kinds of emotions. Moreover, the multimodal embedding space is more discriminative than the unimodal ones, as evidenced by its more compact intra-class and more separated inter-class distributions. Therefore, this comparison result qualitatively verifies the effectiveness of multimodal fusion for audio-visual emotion recognition.
@@ 5 Conclusion
It should be noted that, due to limited computational resources, we cannot afford the pre-training of larger models with more training time and data. Therefore, we plan to explore these issues in future work and encourage other researchers to conduct follow-up studies. Finally, it is also interesting to apply HiCMAE to other audio-visual tasks (e.g., active speaker detection, deepfake detection, and talking face generation).
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) ( No.62276259, No.62201572, No.U21B2010, No.62271083, No.62306316).

### v2

@@ 4.3.3 Results on MSP-IMPROV, RAVDESS, and IEMOCAP
The audio-visual results of MSP-IMPROV [68] are shown in Table 8. We observe that, when compared with the state-of-the-art AW-HuBERT, HiCMAE-B achieves similar UAR but shows a large improvement (+3.15%) in terms of WAR. Two smaller versions of HiCMAE have lower UAR. However, their WAR is still higher than AW-HuBERT.
@@
LR+eGeMAPS [99, 32] A - - 50.30 -
LR+wav2vec [99, 47] A - - 68.80 -
Wav2Vec2.0 [47] A 95 18 73.44 74.38
HuBERT [48] A 95 18 74.15 74.37
WavLM-Plus [71] A 95 18 75.28 75.36
@@
VO-LSTM [100] $\times$ V - - - 60.50
3D ResNeXt-50 [95] $\times$ V 26 - - 62.99
SVFAP [49] V 78 44 75.15 75.01
MAE-DFER [50] V 85 50 75.91 75.56
@@
AV-LSTM [100] $\times$ A+V - - - 65.80
AV-Gating [100] $\times$ A+V - - - 67.70
MCBP [107] $\times$ A+V 51 - - 71.32
MMTM [108] $\times$ A+V 32 - - 73.12
MSAF [95] $\times$ A+V 26 - - 74.86
ERANNs [109] $\times$ A+V - - - 74.80
CFN-SR [96] $\times$ A+V 26 - - 75.76
MATER [104] $\times$ A+V - - - 76.30
MulT [51] $\times$ A+V - - - 76.60
AVT [110] $\times$ A+V - - - 79.20
@@
Attn. Pooling [20]
@@
FBANK [98] $\times$ A - - - 51.52
AV-HuBERT [106] A 90 - - 58.54
RepLAI [111] A 5 - - 57.53
AVBERT [112] A 10 - - 60.94
MAViL [112] A 86 - - 59.46
Wav2vec 2.0 [47] A 95 18 69.88 67.32
HuBERT [48] A 95 18 68.33 66.34
WavLM-Plus [71] A 95 18 68.64 67.12
@@
HoG [34] $\times$ V - - - 35.83
AV-HuBERT [106] V 103 - - 26.59
RepLAI [111] V 15 - - 40.72
AVBERT [112] V 37 - - 45.80
MAViL [112] V 87 - - 43.03
@@
AV-HuBERT [106] A+V 103 - - 46.45
AVBERT [112] A+V 43 - - 61.87
MAViL [112] A+V 187 - - 54.94
@@
In Table 9, we compare the proposed method with state-of-the-art methods on RAVDESS [69]. It can be seen that HiCMAE brings large gains over the previous best method. In specific, our HiCMAE-B achieves a gain of +3.19% WAR over VQ-MAE-AV+Query2Emo [20]. HiCMAE-T still outperforms it by +1.31% WAR, while using 33% fewer parameters. The unimodal results are less satisfactory on this dataset, probably due to the much smaller model size.
The performance comparison on IEMOCAP [70] is presented in Table 10. Most baseline results are from AV-SUPERB [98]. We find that HiCMAE achieves substantial improvement over state-of-the-art generic audio-visual representation learner (e.g., MAViL [19] and AVBERT [112]) in the audio-visual setting. For example, our HiCMAE-S surpasses AVBERT by +5.62% WAR while having a similar model size. HiCMAE-T outperforms MAViL which is also built upon masked autoencoders by +11.68% WAR while being 9$\times$ smaller, indicating the importance of reducing domain shift and the benefit of hierarchical feature learning to improve the quality of the learned representations. When evaluating visual-only performance, our method also demonstrates great success. However, the audio-only results are inferior to state-of-the-art large pre-trained speech models, leaving much room for future performance improvement.
@@ 4.4.1 Datasets
Werewolf-XL [23] is an audio-visual database for studying spontaneous emotions during competitive group interactions in Werewolf games. It contains a total of about 15 hours of audio-visual recordings. In this paper, we use 14,632 speakers’ samples with dimensional annotations (i.e., arousal, valence, and dominance) and conduct subject-independent 5-fold cross-validation for model evaluation.
AVCAffe [24] is a large-scale audio-visual affect dataset simulating remote work scenarios. It consists of a total of 108 hours of videos (more than 58,000 video clips) along with self-reported labels for cognitive load and affect (i.e., arousal, and valence). Note that the arousal and valence scores are given on a scale of 1-4 and we follow the original paper to formulate their prediction as a classification task instead of a regression one. This dataset provides an official split (86 subjects for training and 20 subjects for test) for model evaluation.
@@
eGeMAPS [32] $\times$ A 23.45 28.16 8.08 6.86 31.15 34.03
VGGish [41] $\times$ A 22.88 30.50 5.69 3.91 29.59 31.57
@@
HOG [34] $\times$ V 20.82 14.43 52.54 34.56 24.76 16.90
VGGFace [113] $\times$ V 7.24 4.61 62.96 60.38 14.30 8.20
SVFAP [49] V 23.51 18.96 67.11 64.27 34.61 29.69
@@
Zhang et al. [23] $\times$ A+V 16.41 27.70 63.14 62.34 35.40 38.40
@@ 4.4.2 Results
In Table 11, we compare HiCMAE with state-of-the-art methods on Werewolf-XL [23]. It can be seen that our method outperforms baselines by large margins in terms of Pearson correlation coefficient (PCC) and concordance correlation coefficient (CCC). In the audio-visual setting, for example, HiCAME-B surpasses the previous best by +17% PCC and +4% CCC in arousal and +6% PCC and +4% CCC in valence. In the unimodal setting, HiCMAE also achieves competitive or even better results. Specifically, for the video modality, HiCMAE-B has similar performance with the state-of-the-art SVFAP while having significantly fewer parameters (32M versus 78M) and FLOPs (32G versus 44G). For the audio modality, HiCMAE brings large performance gains over the best-performing baselines on two evaluation metrics in three emotion dimensions except for CCC in arousal.
The audio-visual performance comparison on AVCAffe [24] is presented in Table 12. We find that HiCMAE-B outperforms the previous best results by +2.68% weighted F1-score in arousal and +3.50% weighted F1-score in valence. As for the smaller model HiCMAE-S, although we see some performance drop, it still beats the baselines moderately in two dimensions while having similar parameters with them. Finally, HiCMAE-T typically lags behind the previous supervised methods, probably due to its too-small model capacity.
@@ 4.5 Ablation Studies
In this section, we conduct in-depth ablation studies to investigate several key design factors in HiCMAE. By default, we present the results of HiCMAE-B on two representative datasets, i.e., MAFW (11-class) and CREMA-D (6-class).
Table 13: Ablation study on three-pronged strategy for hierarchical feature learning. HSP: hierarchical skip connections. HCMCL: hierarchical cross-modal contrastive learning. HFF: hierarchical feature fusion.
@@ 4.5.1 Three-pronged Strategy for Hierarchical Feature Learning
Table 14: Ablation study on loss weight for self-supervised pre-training.
@@ 4.5.4 Information Flow in Cross-modal Fusion Encoder
Figure 5: Different types of information flow in cross-modal fusion encoder.
@@
raw-input 42.30 55.86 84.47 84.43
video-first 42.74 56.06 84.81 84.78
audio-first 42.44 56.00 84.60 84.57
@@
We then investigate the effect of different types of information flow in the cross-modal fusion encoder. We develop three variants of the default information flow in Eq. (5-6) and show their differences in Fig. 5. Specifically, for the raw-input variant, tokens of one modality in each fusion layer always attend to the raw input tokens of the other modality [51], instead of updated tokens from the last layer. For the video-first variant, video tokens first update themselves via audio information from the last fusion layer and then audio tokens attend to the updated video tokens. The audio-first variant is just the reverse of the video-first variant. The ablation results are presented in Table 16. We observe that the model performance is not sensitive to different types of information flow in the cross-modal fusion encoder. Besides, in general, the default information flow works best, followed by the video-first and audio-first variants, and finally the raw-input variant.
@@ 4.5.5 Pre-Training Epochs
(a) Total
(b) Fold 1
(c) Fold 2
(d) Fold 3
(e) Fold 4
(f) Fold 5
Figure 10: Confusion matrices on MAFW (11-class). AN: anger. DI: disgust. FE: fear. HA: happiness. NE: neutral. SA: sadness. SU: surprise. CO: contempt. AX: anxiety. HL: helplessness. DS: disappointment.
@@ 4.6.1 Masked Audio-Visual Reconstruction
We first show the reconstruction ability of HiCMAE under different masking rates. We randomly select video clips of four celebrities from the unseen test set of VoxCeleb2 for evaluation. The results of masked audio-visual reconstruction are shown in Fig. 8. For each video clip, we display the original audio spectrogram and 8 facial frame images in the first row, the masked input in medium level (60% for audio and 75% for video) along with the corresponding reconstructed data in the second and third row, and the highly masked input (80% for audio and 90% for video) along with the corresponding reconstructed data in the last two rows. From the figure, we observe that, although some fine-grained details are lost, the global structures of the audio spectrogram (e.g., harmonics) and visual frames (e.g., smiles) are well recovered under both medium and high masking rates. These satisfactory reconstruction results indicate that HiCMAE is capable of fully exploiting limited intra-modal and cross-modal contextual information to infer the missing audio-visual data. We believe this capability learned during self-supervised pre-training has laid the foundation for its superior downstream fine-tuning performance.
@@ 4.6.2 Embedding Space
In this part, we utilize t-SNE [114] to visualize the learned feature embedding space of HiCMAE on CREMA-D (6-class). To demonstrate the effect of audio-visual fusion, we present both unimodal and multimodal embedding space. The results are shown in Fig. 9. Each row presents unimodal or audio-visual embedding space and each column denotes one of five folds on this dataset. From the figure, we find that both audio-only and video-only models have learned good embedding space for distinguishing different kinds of emotions. Moreover, the multimodal embedding space is more discriminative than the unimodal ones, as evidenced by its more compact intra-class and more separated inter-class distributions. Therefore, this comparison result qualitatively verifies the effectiveness of multimodal fusion for audio-visual emotion recognition.
4.7 Error Analysis
We present the confusion matrices on MAFW (11-class) in Fig. 10. The aggregated confusion matrix across five folds is shown in Fig. 10a and the confusion matrix of each fold is shown in Fig. 10b-10f. As can be seen, although achieving significant improvement over previous methods as stated in Section 4.2.2, the performance of HiCMAE on several rare emotions (such as contempt, disappointment, and disgust) is not satisfactory. This is mainly attributed to the imbalanced emotion distribution in real-world scenarios. The imbalanced learning strategies (e.g., resampling techniques and cost-sensitive learning) can be utilized to address this problem and we leave it in future work. We also notice that, among all emotions, neutral emotion is the one most easily confused with other emotions. We believe that the reason is that the boundary between neutral emotion and other emotions is typically less clear than those between other emotion combinations (e.g., happiness and anger). Besides, we observe that fear is often misclassified as surprise. This is consistent with our expectation as their difference in terms of facial expressions is very subtle (both featuring wide-open eyes and raised eyebrows).
@@ 5 Conclusion
It should be noted that, due to limited computational resources, we cannot afford the pre-training of larger models with more training time and data. Therefore, we plan to tackle these issues in future work and also encourage other researchers to conduct follow-up studies. Besides, although HiCMAE has achieved exceptional performance on many datasets, this paper has conducted limited exploration into its internal mechanisms. Therefore, it is worth investigating and enhancing the interpretability of HiCMAE in future work. Possible interpretation tools include attention visualization [73], relevancy map [115], and Grad-CAM [116]. Finally, it is also interesting to apply HiCMAE to other audio-visual tasks (e.g., active speaker detection, deepfake detection, and talking face generation).
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) (No.61831022, No.62276259, No.62201572, No.U21B2010), Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park (No.Z211100004821013), Open Research Projects of Zhejiang Lab (No.2021KH0AB06), and CCF-Baidu Open Fund (No.OF2022025).